# Towards Characterizing Subjectivity of Individuals through Modeling Value Conflicts and Trade-offs

### v1

@@ 5 Experiments
In addition to our proposed framework using LLMs for reasoning, we implement trainable language models varying in structures and subjective ground representations for better comparison. Macro F1 score is used as an evaluation metric, as the label distribution is unbalanced.
@@
The details of the implementation are described in Appendix B, and the codes will be released for future reproduction.
@@ 5.1 Baseline Models
We implement three encoder-only models, DistilBERT (Sanh et al., 2019), RoBERTa (Liu et al., 2019), and DeBERTa-v3 (He et al., 2021), where the input for each model is a text description of the situation and the output is the redditor’s acceptability judgments, 0 or 1.
Moral Judgment Prediction
Model F1-overall F1-skewed
Encoder-only Models
DistilBERT 68.71 52.78
RoBERTa $+$ LoRA 70.48 53.87
DeBERTa-v3 70.94 59.32
Encoder-Decoder Models
BART 68.50 53.55
FLAN-T5 65.44 47.79
Hierarchical Models
BART $+$ DeBERTa 66.67 55.75
FLAN-T5 $+$ DeBERTa 66.86 54.38
gpt-4o as Classifier
Random author’s Comm 72.17 64.53
Same author, random Comm 74.99 73.26
5-shot with Title $+$ Judgment 72.85 66.95
5-shot with Comment 76.78 72.89
5-shot with Diff Judgment 77.43 74.44
SolAr: gpt-4o Clf, Selection w/ Abstraction
Schwartz’s Basic Values 75.81 72.95
[…]
Table 2: Macro F1 scores averaging all redditors (-overall) and two redditors with extremely skewed label distributions (-skewed).
@@
We also fine-tune encoder-decoder language models, mainly BART (Lewis, 2019) and FLAN-T5 (Chung et al., 2024), to learn more about each redditor’s subjectivity. The models get the same input, but their objective is to generate the redditors’ likely reactions (i.e., comments and judgments). While encoder-only models only observe each redditor’s binary judgment patterns with respect to the situations, the models fine-tuned with this approach have access to the texts that are authored by the redditors.
Finally, we implement hierarchical models that combine language generation abilities of encoder-decoder models and classification abilities of encoder-only models. In this setup, encoder-decoder models generate each redditor’s likely reactions without judgments, and a separate encoder-only model learns likely judgments from the generated comments.
Since the number of training instances for each redditor might not be sufficient, we try techniques that help fine-tuning the models with smaller number of instances such as LoRA (Hu et al., 2021). We report the best combinations resulting for each model structure in Table 2.
5.2 Subjective Ground Selection
As described in 4.1, we represent the subjective ground of an individual in four different ways: using comments from similar situations, comments from similar situations while presenting different judgments, abstract values generated from comments using fixed value definitions (i.e., Schwartz’s theory of basic human values), and abstract value trade-offs generated by LLMs. We choose the top five items as few-shot examples for each instance and prompt the gpt-4o model to get the final judgment prediction.
In order to compare the effectiveness of subjective ground retrieval, we prompt the same gpt-4o model with randomized few-shot examples in two ways, random redditors’ comments, and the target redditor’s random comments. Random redditors’ comments are neither relevant to the test situations nor represent subjective preferences of the target redditor, while the target redditor’s random comments provide what the target redditor has commented on other irrelevant situations.
@@ 6.1 Inference Performance
[…]
@@
We also observe that as the model becomes more complex and parameterized, the inference performance worsens—the F1 scores of encoder-decoder models are worse than encoder-only models, and hierarchical models are even worse than that. This implies that the nature of data scarcity in subjectivity analysis also makes it more difficult to fine-tune a specific individual’s perspectives to characterize their subjectivity.
@@
Selecting the most relevant subjective ground of redditors given a test situation works generally well, answering one of our research questions, “Can off-the-shelf LLMs account for individual’s subjective preference?”. As RAG-based inference does not require training, the performance of F1-skewed becomes more similar to the overall F1 scores.
Among different selection methods, using vector representations of value trade-offs work the best. Mapping comments to a fixed set of value systems suggested by Schwartz (1992) performs worse than using redditors’ comments as is. This is potentially because ten fixed set of values are too abstract to map diverse perspectives of individuals on social situations. Selecting few-shot examples with enforcing different moral judgments perform the second best, implying that it is important to provide more diverse examples of redditors’ subjective preferences in order to better ground LLMs.
Moral Judgment Prediction - Controversial Situations
Model F1-overall F1-skewed
Encoder-only Models
DistilBERT 58.15 42.04
RoBERTa 59.95 49.35
DeBERTa-v3 56.32 56.93
gpt-4o as Classifier
Random Author’s Comm 52.65 56.62
Same author, random Comm 55.40 59.61
5-shot with Title $+$ Judgment 56.45 59.27
5-shot with Comment 55.66 59.27
5-shot with Diff Judgment 57.45 62.30
SolAr: gpt-4o Clf, Selection w/ Abstraction
Schwartz’s Basic Values 55.00 61.39
[…]
Table 3: Macro F1 scores on controversial situations. RAG-based methods, except for our proposed framework, perform worse than fine-tuned encoder-only classifiers.
Figure 2: Heat map of value trade-offs among the top 8 redditors. There are some pairs of conflicting values where the majority of redditors show similar preference (e.g. tradeoff_C), yet in most cases their value trade-offs work differently.
6.2 Performance on Controversial Situations
One of the surprising results from Table 2 is that prompting with random comments still perform reasonably well. We hypothesize that LLMs tend to judge the situations on their own when the few-shot examples no longer provide relevant subjective choices, and their judgment patterns would look similar to humans’ if the situation is less controversial.
[…]
6.3 Do people show diversities in value trade-offs?
@@
A key research hypothesis of this study is that each redditor actively participating in r/AmITheAsshole exhibits distinct subjectivity patterns, making the analysis of individual perspectives a fundamentally different NLP challenge compared to social commonsense reasoning. We validate this hypothesis by visualizing the value trade-off patterns of the top redditors.
[…]
@@
The heat map tells us that there are some values commonly cherished by most of the redditors, yet they show different priorities and trade-offs for other values. For instance, when looking at the third column (tradeoff_C), almost all redditors prioritize “Effective Communication and Relationships” over “Emotional Influence and Accountability”. Redditors show various patterns when it comes to “Personal Autonomy and Freedom” (tradeoff_F to tradeoff_I), and especially for redditor 8, this value is prioritized in most cases, unless it conflicts with “Respect and Compliance with Authority”. This visualization not only suggests that active redditors in r/AmITheAsshole show distinct subjectivity, but explains their value preferences and judgments.
Figure 3: Euclidean distance of pairs of value representations with respect to their corresponding Pearson correlation coefficient computed with value win rates.
6.4 Do people show consistency in value preferences?
The idea of considering past comments as subjective ground and using them to infer redditors’ likely judgments on unseen situations is logically valid only if redditors exhibit consistent judgment patterns. For instance, if a redditor consistently favors one value, A, over a conflicting value, B, it is reasonable to infer that they would also prefer a similar value, A, assuming their decision-making remains consistent.
To show the consistency of value preferences, we plot the distance between pairs of value representations with respect to their corresponding correlation coefficients. We compute Pearson correlation coefficient using value win rates; when a value has a high win rate, another value that has correlation coefficients close to 1 also has a high win rate. Thus, more similar values should have correlation coefficients near 1, while increasing distance between values is expected to result in coefficients approaching -1.
Figure 3 illustrates the pairwise Euclidean distance of value representations on the x-axis and their Pearson correlation coefficients on the y-axis. The regression plot indicates a clear trend: as the similarity between two values increases, their correlation becomes more positive, whereas further distances correspond to more negative correlations. This analysis suggests that redditors show consistent value preferences, thereby validating our approach of using one’s past comments to infer their likely judgments.
@@ 7 Related Studies
The closest neighbor of this research is individual subjectivity analysis. Lee and Goldwasser (2022) analyzes the same community, r/AmITheAsshole and learned subjective preference of individuals by computing attention weights between rules-of-thumb and situations. In the political framing and agenda-setting domain, Roy and Goldwasser (2021) utilized Moral Foundations theory to characterize real world politicians varying in topics. Representing individual-level perspectives is also done in debate setting; Li et al. (2018) used graph embeddings to associate opnions and individuals together.
More recently, researchers utilized LLMs to ground their generations to a desired persona or personality. This persona-based control is done in many domains including basic question answering (Zheng et al., 2024a), interactive simulacra (Park et al., 2023), and solving causal inference and moral dilemma (Nie et al., 2023). Choi and Li (2024) proposed Persona In-Context Learning which uses Bayesian inference to select the optimal set of persona for a given task. Researchers have also tried providing more direct signals in the prompt using demographic information (Durmus et al., 2023).
Another line of related studies is the Retrieval-Augmented Generation. RAG retrieves the most relevant information to mainly fine-tune language models or help off-the-shelf LLMs infer better on many downstream tasks such as QA (Shi et al., 2023; Xu et al., 2023; Wang et al., 2024b), reasoning and language understanding (Yu et al., 2023; Lin et al., 2023; Zhang et al., 2023), text summarization and generation (Guo et al., 2023; Jiang et al., 2023; Yan et al., 2024), etc. Researchers have also studied whether RAG can be applied to more personal use case such as recommendation system (Rajput et al., 2023) and personalized dialog generation (Wang et al., 2023, 2024a). These approaches only consider factual information (e.g. purchase history, where did the person go two days ago) as a personalized aspect, while our research characterizes subjective perspectives and applies RAG for performance improvement.
@@ 8 Conclusion
In this paper, we propose a framework, SolAr, that takes redditors’ past comments into account and characterizes their subjective ground using value abstraction. Empirical results show that LLMs can be efficiently used to account for subjective preference of individuals compared to traditional methods that require fine-tuning. We also show that the performance can be further improved by retrieving the most relevant subjective ground using value trade-offs. Furthermore, SolAr provides an additional explanations about each redditor’s distinct value preference patterns which could later be used to justify LLM inference.
@@
Although it is not feasible to model subjectivity of individuals that is perfectly correct and inclusive, representing it with their past comments is still an over-simplified definition of subjectivity. In the future studies, we plan to incorporate more redditor-related information such as their community membership (i.e., what other subreddits they are actively participating) and activities on other communities to better characterize individuals.
@@
[…]
@@
Appendix A Value Abstraction
A.1 Schwartz’s Theory of Basic Human Values
@@
  • Achievement: “personal success through demonstrating competence according to social standards”
@@
  • Conformity: “restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms”
  • Tradition: “respect, commitment, and acceptance of the customs and ideas that one’s culture or religion provides”
  • Benevolence: “preserving and enhancing the welfare of those with whom one is in frequent personal contact (the ’in-group’)”
  • Universalism: “understanding, appreciation, tolerance, and protection for the welfare of all people and for nature”
In a more generalized view of the values, these values can be grouped into four categories, Openness to Change (self-direction, stimulation), Self-Enhancement (hedonism, achievement, power), Conservation (security, conformity, tradition), and Self-Transcendence (benevolence, universialism). Openness to Change and Conservation contradicts to each other, and Self-Enhancement and Self-Transcendence contradicts to each other as well.
In order to annotate comments to these fixed values, we first prompt gpt-4o model to generate values observed from the comments in general. Then we prompt again, asking “map the value items to the most relevant dimensions that Schwartz has defined in his Theory of Basic Human Values”.
A.2 Clustering Value Conflicts and Trade-offs

### v2

@@ 5 Experiments
In all experiments, we use the macro F1 score as the evaluation metric due to the imbalanced label distribution. For each model, we first calculate the macro F1 score individually for each redditor, then take the unweighted average of these scores across all redditors. This approach ensures that every redditor has equal influence on the final score, regardless of how many instances each contributed. Even if a model performs well on some redditors who are easily predictable, its overall score becomes lower if it fails to learn the subjectivity of other redditors.
Moral Judgment Prediction - Baselines
Model All Redditors Top 8 Redditors
Encoder-only Models
DistilBERT-base-uncased 47.90 $^{\pm\text{0.0053}}$ 68.71 $^{\pm\text{0.0145}}$
RoBERTa-base 46.68 $^{\pm\text{0.0038}}$ 70.48 $^{\pm\text{0.0088}}$
DeBERTa-v3-large 42.45 $^{\pm\text{0.0013}}$ 70.94 $^{\pm\text{0.0024}}$
Encoder-Decoder Models
BART-base 48.90 $^{\pm\text{0.0055}}$ 68.50 $^{\pm\text{0.0150}}$
FLAN-T5-base 46.21 $^{\pm\text{0.0035}}$ 65.44 $^{\pm\text{0.0118}}$
Encoder-Decoder Models; Seq2Seq
BART-base 40.83 $^{\pm\text{0.0012}}$ 66.67 $^{\pm\text{0.0086}}$
FLAN-T5-base 43.01 $^{\pm\text{0.0020}}$ 66.86 $^{\pm\text{0.0087}}$
Table 2: Macro F1 scores of baseline models. The average performance of fine-tuned models over all 100 redditors shows a significance drop compared to the performance of the top 8 redditors who commented over 1,000 times.
Moral Judgment Prediction - RAG-based Models
Retrieval Strategy Input All Redditors Top 50% Redditors Bottom 50% Redditors
All Contro All Contro All Contro
@@
[…]
Schwartz’s Value Comment 79.05 $^{\pm\text{0.003}}$ 81.38 $^{\pm\text{0.006}}$ 78.29 $^{\pm\text{0.005}}$ 80.97 $^{\pm\text{0.004}}$ 79.81 $^{\pm\text{0.002}}$ 81.82 $^{\pm\text{0.007}}$
[…]
Table 3: Performance of different retrieval strategies across author percentiles with respect to data size. “Contro” denotes controversial situations. SolAr, which ensembles different strategies and input types for controversial and non-controversial situations at test time, achieves the best overall performance. The performance boost is more significant for the bottom 50% redditors and in controversial situations.
@@ 5.1 Baseline Models
We implement three encoder-only models, DistilBERT (Sanh et al., 2019), RoBERTa (Liu et al., 2019), and DeBERTa-v3 (He et al., 2021), where the input for each model is a text description of the situation and the output is the redditor’s acceptability judgments, 0 or 1. We denote this approach as Endoer-only Models.
Denoted as Endoer-Decoder Models, we fine-tune encoder-decoder language models, mainly BART (Lewis, 2019) and FLAN-T5 (Chung et al., 2024). We report the experimental results of these models with the same classification setting as the encoder-only models; models get situation text as input and output either 0 or 1.
@@
To learn more about each redditor’s subjectivity, we test another variation. The models get the same input, but their objective is to generate the redditors’ likely reactions (i.e., comments and judgments). While encoder-only models only observe each redditor’s binary judgment patterns with respect to the situations, the models fine-tuned with this approach have access to the comments that are authored by the redditors. This approach is denoted as Endoer-Decoder Models; Seq2Seq. Table 2 illustrates the results of the baseline model. Further details of the baseline model implementation are described in Appendix C.1.
5.2 RAG-based Models
We compute the performance for each redditor separately, using the same test set as the ones used in baseline models. As described in 4.1, we represent the individual subjectivity in a several different ways.
First, we try different retrieval strategies by querying past history data with situations and abstract values that are annotated by LLMs. For querying with abstract values, we report the results of using Schwartz’s values and value trade-offs. After retrieval, we differentiate LLM inputs (i.e. few-shot examples) by adding Schwartz’s values and value trade-offs that are observed from each redditor’s past history.
Our proposed framework, SolAr, ensembles retrieval strategies based on the difficulty of the test instance. When the test situation is given, SolAr computes its difficulty based on how controversial it is—defined as having less than 70% agreement among all redditors’ judgments [^4] This is the threshold that r/AITAFiltered uses to identify controversial situations from r/AmITheAsshole.. For controversial situations, SolAr employs the value-aware retrieval function $R_{val}(\ )$, while it uses the situation-based retrieval function R()R(\ ) for non-controversial cases. In both scenarios, few-shot examples include both the comment and the corresponding value trade-offs.
We use GPT-4.1 (OpenAI, 2025) to predict the acceptability judgments of individuals. Each method retrieves five few-shot examples and prompts the LLM twice to estimate a confidence interval. Table 3 shows the performance of each method. Appendix C.2 explains more detailed experimental settings for RAG-based models.
Figure 2: Average distance between the test situation and retrieved instances for each redditor. $x$-axis is the log-scaled numbers of each redditor’s data size. When the redditor has more instances, it is more likely that they have commented on situations that are more similar to the test situation in the past.
@@ 6.1 Inference Performance
The overall F1 scores of the baseline models in Table 2 show that fine-tuning redditors’ decision patterns with language models does not solve the problem. The models show fairly good performance only for the redditors who have enough amount of training instances, while they fail to learn the subjectivity of redditors who have a small number of instances and highly skewed judgment patterns. We visualize fine-tuned models’ performance with respect to the number of instances and judgment skewness in Appendix D.
@@
We also observe that as the model becomes more complex and parameterized, the inference performance worsens—the F1 scores of encoder-decoder models are worse than encoder-only models, and sequence-to-sequence models are even worse than that. This implies that the nature of data scarcity in subjectivity analysis also makes it more difficult to fine-tune a specific individual’s perspectives to characterize their subjectivity.
@@
On the other hand, RAG-based models that select the most relevant instances to a test situation work generally well. This answers one of our research questions, “Can off-the-shelf LLMs account for individual’s subjective preference?”. As RAG-based inference does not require training, the performance of redditors with less amount of instances (i.e. Bottom 50%) matches the overall F1 scores.
Figure 3: Visualization of subjective preferences. The heatmap shows the most common value conflicts among the 8 most active redditors. Each value trade-offs are represented in a “A > B” form, and each cell’s value means the win rate of value A over value B for a specific redditor. In most cases, redditors’ value trade-off systems work differently.
Different retrieval methods and inputs show distinct advantages. First of all, when we compare different inputs within the standard retrieval strategy (i.e. Situation), adding values improves the performance for redditors who have less amount of instances; using comment and value trade-offs improves the F1 score by 2.69% for the bottom 50% redditors. Figure 2 shows that the average distance between the test situation and the retrieved instances increases for redditors with fewer instances. The performance boost we get for the bottom 50% redditors implies that adding abstract values helps characterize the redditors’ subjectivity when the retrieved instances are topically less similar to the test situation.
Abstract values are also helpful when we use them for retrieval. The performance in controversial situations improves when the most relevant instances are retrieved using either Schwartz’s values or value trade-offs. This suggests that predicting individuals’ subjective preferences benefits from knowing their high-level value system, especially when the test situation is controversial and LLM’s own morality does not align well with humans.
6.2 Diverse Value Trade-offs among Redditors
@@
A key research hypothesis of this study is that each redditor actively participating in r/AmITheAsshole exhibits distinct subjectivity patterns, making the analysis of individual perspectives a fundamentally different NLP challenge compared to social commonsense reasoning. We validate this hypothesis by visualizing the value trade-off patterns of the 8 most active redditors.
Figure 3 displays the 8 most common value trade-offs of the most active redditors. Each cell in the heatmap refers to win rates. For example, $0.9$ in the cell in row 1, column 1 means that among all situations where a value “Boundary Awareness” conflicts to “Cultural Expectations”, redditor 1 chooses the former over the latter 90% of the times. For the same situations, redditor 3, who has a $0.45$ win rate, chooses “Boundary Awareness” only 45% of the time, implying they prefer “Cultural Expectations” slightly more.
@@
The heatmap tells us that there are some values commonly cherished by most of the redditors. The bottom row of the heatmap, for instance, shows that all redditors prefer “Personal Autonomy” over “Parenting Decision-Making” in most cases. At the same time, redditors show diverse patterns when it comes to “Personal Autonomy” conflicts with other values. For redditor 8, specifically, this value is highly prioritized in most cases, unless it conflicts with “Compliance with Authority”. This visualization not only suggests that active redditors in r/AmITheAsshole show distinct subjectivity, but explains their value preferences and judgments.
@@ 7 Related Studies
The closest neighbor of this research is individual subjectivity analysis. Lee and Goldwasser (2022) analyzes the same community, r/AmITheAsshole and learned subjective preference of individuals by computing attention weights between situations and social norms. Plepi et al. (2022) used the authorship information with text embeddings and perform the YTA/NTA classification of the redditors in r/AmITheAsshole. In the political framing and agenda-setting domain, Roy and Goldwasser (2021) utilized Moral Foundations theory to characterize real world politicians varying in topics. More recently, researchers utilized LLMs to ground their generations to a desired persona or personality in many domains, including question answering (Zheng et al., 2024a), interactive simulacra (Park et al., 2023), and solving causal inference and moral dilemma (Nie et al., 2023). Choi and Li (2024) proposed Persona In-Context Learning which uses Bayesian inference to select the optimal set of persona for a given task. Researchers have also tried providing more direct signals in the prompt using demographic information (Durmus et al., 2023).
Incorporating value theories into LLMs is an active research area these days. Van Der Meer et al. (2023) analyzed (dis-)agreements between the users by identifying value profiles with Schwartz’s value theories. Sorensen et al. (2025) used LLMs as proxy humans with different attributes (e.g. demographic traits, value profiles), and analyzed their distinct behaviors on controversial issues. Bhatia et al. (2025) used LLMs to characterize individuals’ choices by generating descriptions of benefits and costs of their choices. Sorensen et al. (2024) used concrete examples (i.e. text document) and prompted LLMs to generate corresponding values, rights, and duties. Ye et al. (2025) parsed free-form input text into various perceptions and the corresponding Schwartz values by fine-tuning LLMs.
The main differences between our approach and the existing studies in subjectivity analysis and human value analyses with LLMs are; (1) we analyze actual redditors in real world rather than using LLMs as proxy humans, (2) our approach applies more fine-grained buckets for subjectivity, and (3) the novel value trade-off analysis provides additional explanation for individuals’ choices.
Another line of related studies is the Retrieval-Augmented Generation. RAG retrieves the most relevant information to mainly fine-tune language models or help off-the-shelf LLMs infer better on many downstream tasks such as QA (Shi et al., 2023; Xu et al., 2023; Wang et al., 2024b), reasoning and language understanding (Yu et al., 2023; Lin et al., 2023; Zhang et al., 2023), text summarization and generation (Guo et al., 2023; Jiang et al., 2023; Yan et al., 2024), etc. Researchers have also studied whether RAG can be applied to more personal use cases such as recommendation system (Rajput et al., 2023) and personalized dialog generation (Wang et al., 2023, 2024a). These approaches only consider factual information (e.g. purchase history, where the person went two days ago) as a personalized aspect, while our research characterizes subjective perspectives and applies RAG for performance improvement.
@@ 8 Conclusion
In this paper, we propose a framework, SolAr, that takes redditors’ past comments into account and characterizes their subjective ground using value abstraction. Empirical results show that LLMs can be efficiently used to account for the subjective preference of individuals compared to traditional methods that require fine-tuning. We also show that the performance, especially for redditors with small data sizes and in controversial situations, is improved by retrieving the most relevant instances using abstract values. Furthermore, SolAr provides additional explanations about each redditor’s distinct value preference patterns which could later be used to justify LLM inference.
@@
Although it is not feasible to model the subjectivity of individuals that is perfectly correct and inclusive, representing it with their past comments is still an oversimplified definition of subjectivity. First, past comments in the same community might not cover all aspects of the subjectivity. In future studies, we plan to incorporate more redditor-related information such as their community membership (i.e., what other subreddits they are actively participating in) and activities in other communities to better characterize individuals. Another oversimplification is that it assumes the redditors would judge the situations consistently over time. It would be an interesting direction to analyze whether there are value shifts over time for the redditors.
@@
In terms of downstream tasks and the model’s performance, our proposed framework shows sub-optimal performance in predicting the correct judgment. Although our framework shows higher improvements when considering controversial situations, our goal is to make LLMs perform well not only in controversial situations but also in other situations.
@@
Acknowledgements
This project was partially funded by NSF IIS-2048001 and DARPA CCU program. The views are the authors’ and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government.
@@ Acknowledgements
Appendix A Data Distribution
Figure 4: PCA analysis of situation representations of all r/AmITheAsshole posts and truncated posts we use for training and inference. Truncated situations show similar distributions to the population distribution.
As described in 3.1, we crawl all posts in r/AmITheAsshole from November 2014 to June 2023 and filter out threaded comments. As a result, we have around 217K unique situations with 891K instances in total. Fine-tuning models and performing LLM inference on this massive dataset takes up too many resources. Thus we truncate the dataset. We first identify the most active redditors (i.e. redditors who commented the most). Then we identify the redditors who commented more than a certain threshold [^5] We set this as 2,000. We get 8 redditors from this process. After this step, we filter out situations that the 8 most active redditors did not leave comments. As a result, 1.7K unique situations are left.
Figure 5: Redditors’ judgment patterns with respect to the redditor counts in the truncated data (top), and in r/AmITheAsshole (bottom).
To ensure that the truncated situations and all redditors who commented on these situations are not too different from the population distribution, we first visualize situation representations in a 2-D embedding space using Principal Component Analysis. Figure 4 visualizes the vector representations of all 217K situations in r/AmITheAsshole (blue dots), and vectors of 1.7K situations in the truncated dataset (orange crosses). The plot implies that the situations in the truncated dataset are not deviated from the population distribution.
We compare the statistics of redditors in r/AmITheAsshole and the truncated dataset. Figure 5 shows the redditors’ judgment patterns (i.e. acceptable or not acceptable) in the truncated data and in all r/AmITheAsshole. A judgment pattern of 1 means that the redditor judged all situations as “acceptable”, and 0 means they judged all as “not acceptable”. When the value is 0.5, the redditor has a perfectly balanced judgment history. In the truncated data, redditors tend to judge more situations as “acceptable”, and the trend is the same in the r/AmITheAsshole. This shows that the redditors in the truncated dataset do not diverge from the overall population.
Appendix B Value Abstraction
B.1 Schwartz’s Theory of Basic Human Values
@@
  • Achievement: “personal success through demonstrating competence”
@@
  • Conformity: “restraint of actions likely to upset or harm others and violate social norms”
  • Tradition: “respect of the customs and ideas that one’s culture or religion provides”
  • Benevolence: “preserving and enhancing the welfare of those within-group)”
  • Universalism: “protection for the welfare of all people and for nature”
In order to annotate situations and comments to these fixed values, we prompt gpt-4o-0806 model to generate values observed from the text. Below is an example prompt:
Prompt for Annotating Schwartz’s Values You will be given a social situation and a Redditor’s comment. Your job is to find the top two most salient Schwartz’s Basic Human Values that are observed in the situation and in the comment. Follow the format below without generating any other explanation.
Input Format:
[Situation] situation_description
[Comment] redditor_comment
Output Format:
[Situation] Schwartz_value_A, Schwartz_value_B
[Comment] Schwartz_value_C, Schwartz_value_D
[Situation] Not babysitting my niece
[Comment] When you talk to your brother ..
B.2 Clustering Value Conflicts and Trade-offs
Similar to annotating Schwartz’s values, we prompt gpt-4o-0806 model and generate value conflicts and trade-offs. Below is an example prompt:
Prompt for Annotating Value Conflicts You will be given a situation. Your task is to understand the situation, and tell me what kind of moral values are conflicting.
[Situation] Not babysitting my niece …
Tell me what kind of moral values are conflicting in the situation. Note that the two conflicting values can’t be chosen at the same time (i.e. trade-off), and people’s behaviors and attitudes will be different based on which value they choose.
Highlight one or more core conflicting values that describe the situation the best. Values should be generic; proper nouns or pronouns should not be included. Write values in phrases, not in sentences, and each value should have at least 3 words. Your response should follow the format:
"A vs. B",
"C vs. D",
After the conflicting values are annotated for the situations, we prompt the GPT model again to analyze the comments from each redditor. Below is a prompt:
Prompt for Annotating Value Trade-offs You will be given a [Situation] and a list of [Conflicting Values] observed in the situation. There’s Person X who leaves a [Comment].
[Task] Determine which items in the [Conflicting Values] are mostly related to Person X’s [Comment]. You may select only one of them, or select multiple items. Generate conflicting values observed in the comment ONLY IF none of the [Conflicting Values] are related to the [Comment].
Determine which value is more important to Person X. For instance, "A vs. B" is a chosen conflicting value and Person X thinks A is more important than B in the comment, the answer should be "A > B".
Your response should follow the format:
"A > B",
"D > C",
[Situation] Not babysitting my niece
[Conflicting Values] Autonomy vs. Family
[Comment] When you talk to your brother ..